# Redshift Drift in Relativistic N-Body Simulations

### v1

@@ Abstract
The cosmological redshift drift promises to be the first observable directly measuring the cosmic expansion rate and should be detectable with upcoming surveys by the Square Kilometre Array and the Extremely Large Telescope. To prepare for these upcoming measurements we study the redshift drift in detail using the relativistic N-body code gevolution focusing on inhomogeneity-induced fluctuations. Using a ray-tracer, we calculate the redshift drift directly from the light cone at two different time steps. To investigate observer-dependent biases we consider 10 different observers. We find that inhomogeneity-induced fluctuations in the redshift drift can in extreme cases be of the same order as the cosmic signal for $z\lesssim 0.15$. By comparing our results to first-order perturbation theory, we find that the extreme outliers are due to peculiar motion in over-densities and can be described by first-order perturbation theory to percent precision. We calculate angular power spectra that fit very well with our theoretical predictions based on perturbation theory at linear scales and show a surprisingly large non-linear signal. This shows that redshift drift not only has the power to measure the background expansion, but could also deliver information about the velocity and acceleration fields in clusters.
@@ 1 Introduction
One of the primary tasks in cosmology is to measure the cosmic expansion rate. Such measurements constrain the spacetime geometry, the energy content of our universe, and the theory of gravity describing its evolution. Measurements of the expansion rate through supernovae of type Ia gave us the first hint that our universe is expanding in an accelerated manner, which required us to include dark energy in our model of the Universe [1; 2]. However, so far, all of our measurements of the expansion rate have been indirect. Measurements of supernovae for example constrain the distance-redshift relation and therewith an integrated measure of the expansion rate. A direct measurement of the cosmic expansion rate would be a huge step forward in constraining and understanding our universes’ evolution and the phenomenology of dark energy.
Such a direct measurement of the cosmic expansion is for the first time becoming feasible with the current generation of telescopes. Telescopes such as the Square Kilometre Array (SKA) and the Extremely Large Telescope (ELT) are predicted to be able to measure the real-time variation of the cosmological redshift [3; 4; 5], the so-called redshift drift [6; 7]. The redshift drift is the variation of the redshift, $z$, of a given source observed at two different times $\delta t_{\mathrm{o}}$ apart, i.e.
@@
The real universe is, however, only homogeneous and isotropic, i.e. described by an FLRW metric, if viewed on the largest scales. On smaller scales, the Universe is full of structures. These structures affect any cosmological observables, including the redshift drift. Indeed, the effects of structures have been identified as one of the main obstacles for obtaining unbiased redshift drift measurements [3]. While it was estimated in [3], that the noise will be subdominant to the signal expected for observations with the ELT, unless based on observations of a small number of objects with large peculiar velocity and acceleration, and observations with the SKA will average over billions of galaxies [4], meaning that peculiar motion should only introduce random noise and not bias the result. It remains important to precisely quantify the size of the noise in a realistic simulation. This is both important in order to confirm that the measurements will indeed be unbiased and do not suffer from observer-dependent effects. Furthermore, in addition to being a source of noise in the background signal, the peculiar motion of objects might provide a useful signal in itself. For example [8] suggested that the peculiar acceleration of objects in nearby clusters could be used to map out their gravitational potential. It is therefore interesting to study whether such signals can be extracted from redshift drift measurements. The time variation of the redshift, in particular due to peculiar motion, could also be important for future gravitational wave detectors such as LISA or DECIGO [9; 10].
The fluctuations in the redshift drift have already been studied in several cosmological simulations. In [11], the redshift drift was studied using the Newtonian N-Body code GADGET-2 [12], while [13] studied the redshift drift in fully relativistic simulations of structure formation run with the Einstein Toolkit [14; 15]. Both papers considered an Einstein-de Sitter universe and calculated the redshift drift using the general expansion introduced in [16] keeping leading-order terms. Furthermore, both studies used quite low resolution [^2] The resolution of the actual simulation used in [11] is not particularly low with $512^{3}$ particles on a 512 Mpc/h grid. However, redshift drift computations were produced only after smoothing the data onto the grid. which may be important when assessing the size of the effects peculiar motion has on redshift drift fluctuations. In [17], the redshift drift fluctuations were studied using GADGET-4 [18], calculating the redshift drift from perturbation theory on the spatial background hypersurfaces/snapshots of the simulation. This study was the first to consider redshift drift fluctuations in a simulation based on a $\Lambda$CDM background, but the study only estimated the redshift drift signal on constant coordinate-time snapshots and thus at background redshifts rather than using the fluctuating redshift. In addition, the study compared to linear perturbation theory and found a striking difference between perturbative predictions and simulation results.
The goal of this paper is to address the shortcomings of these earlier studies by computing the exact redshift drift in a $\Lambda$CDM-based relativistic N-body simulation with high resolution, with the aim of understanding the relationship between simulation and perturbative results. To do this, we use the relativistic N-body code gevolution [19].
@@
The paper is structured as follows. We start in Sect. 2 by introducing our simulation setup and explaining how we calculate the redshift drift. Next, we derive the redshift drift and the redshift drift fluctuation power spectrum in a perturbed FLRW universe in Sect. 3. In Sect. 4 we present the results of our simulation and compare them with the results from Sect. 3 and other results from the literature, before concluding in Sect. 5.
@@ 2 Simulation Setup
The results presented in this paper are produced with the relativistic N-body code gevolution [^3] https://github.com/gevolution-code[19]. gevolution employs a particle-mesh approach and implements the perturbed FLRW metric in the Poisson gauge
@@
where $\tau$ is the conformal time, $a$ the background scale factor and $x^{i}$ are co-moving coordinates. Here, $\Psi$, $\Phi$ are the two scalar perturbations, known as the Bardeen potentials, and $B_{i}$, $h_{ij}$ are the vector and tensor perturbations. The metric components are stored on a regular Cartesian grid and evolved according to the field equations of general relativity expanded in a weak field expansion. For the exact details of the expansion and an overview of the terms that are kept, the reader is referred to [20]. The positions and velocities of the particles are updated by solving the geodesic equation.
In addition to the metric on the Cartesian grid and full particle positions and velocities, gevolution can output particle positions and metric fluctuations on background light cones for an observer at a given position and observation time. In this case, the metric is stored on a HEALPix [21] grid.
One can post-process the output with a raytracer to compute the redshift and position of the individual particles on the sky. We use the raytracer first introduced in [22]. The raytracer uses an iterative procedure, shooting light rays from the observer positions towards the source, solving the geodesic equation along the ray using the perturbed metric, though only retaining the scalar perturbations. If the ray misses the source, the initial shooting angle is adjusted for the next pass until the correct geodesic is calculated.
@@
Our simulation uses a flat $\Lambda$CDM model as background, with initial conditions created using CLASS [^5] http://class-code.net [23], using the following cosmological parameters $h=H_{0}/(100\,\mathrm{km}\,\mathrm{s}^{-1}\mathrm{Mpc}^{-1})=0.67556,\;\Omega_{c}h^{2}=0.12038,\;\Omega_{b}h^{2}=0.022032,\;A_{s}=2.215\times 10^{-9}$ (at $0.05\,\mathrm{Mpc}^{-1}$) and $n_{s}=0.9619$, which are the standard parameters for gevolution and very close to standard Planck parameters [24].
@@
3 Theoretical Predictions
The redshift drift in a perturbed FLRW universe was first discussed in [25; 26; 27] and later derived in [28] in a gauge-independent manner. In this paper, we will stick closest to the result by [27], but extend it by a term that appears when observing on the light cone. We repeat the main steps of the calculation to make this paper self-contained and introduce our notation. We will only consider the dominant corrections to the redshift and redshift drift which come from the peculiar motion and not the metric fluctuations. In the derivations below we therefore assume that the wave vector is that of an FLRW spacetime and that the only redshift corrections from inhomogeneities are those from the peculiar motion of the source and observer which will be taken into account at the linear level. We thus assume that the light paths including the null tangent vector are in accordance with an FLRW spacetime with line-element
@@ 3.1 Redshift Drift
Within the geometric optics approximation the wave vector has to be $k^{\mu}=\omega\;(1,\hat{k}^{i})$ [29], where $\omega$ is the angular frequency measured by a local Lorentz observer and $\hat{k}$ is the unit direction vector pointing in the direction of propagation of the wave. Considering an observer with 4-velocity $u^{\mu}=(1,u^{i})$ one finds
@@
Since $\mathbb{n}$ is constant along the geodesic, we are free to choose it at any point along the light ray. Since we further replaced $\mathbb{n}=-\mathbb{e}$ we are now free to choose $\mathbb{e}$ at any point. We will always choose $\mathbb{e}_{\mathrm{o}}$ later, but simply keep writing $\mathbb{e}$ for ease of notation. The result (19) now agrees with the result in [27], but is not yet the one we observe on the light cone. The second effect we have to account for comes about from the following: So far all quantities are evaluated at the background time $t$, related to hypersurfaces of constant background redshift $\bar{z}$. Observationally we do not have access to $\bar{z}$, only to the observed redshift $z=\bar{z}+\Delta z$. For any function we can relate
@@
[30; 28]. Within perturbation theory, such corrections need only be applied to background quantities, i.e. in our case only $H_{\mathrm{s}}$. We have
@@
which will come in handy when deriving the angular power spectrum in the next subsection. Eq. (23) agrees with the expression obtained in [28], except for the observer-dependent term that was neglected there, and minus all terms arising from metric fluctuations, which we neglect here.
@@ 3.2 Angular Power Spectrum
We note that the power spectrum of redshift drift fluctuations has already been calculated to first order in perturbation theory in [28] in a fully gauge-independent manner, accounting not only for the velocity perturbations but also for metric perturbations. Since we choose a different convention for $\Delta\delta z$ ([28] neglect the $H_{0}$ term, which we retain), we present our own derivation here.
@@
$a_{lm}^{\Delta\delta z}(z)=\frac{1}{(2\pi)^{3}}\int\mathrm{d}\Omega_{\mathbb{e}}Y_{lm}^{*}(\mathbb{e})\int\mathrm{d}^{3}k\;\Delta\delta z(\mathbb{k},z)\,\mathrm{e}^{-ir_{\mathrm{s}}\mathbb{k}\cdot\mathbb{e}}\;,$ (29)
where $r_{\mathrm{s}}$ is the co-moving distance to our source. Introducing the velocity potential $v$ such that $\mathbb{v}=\mathbb{\nabla}v$, we simply have $\mathbb{v}=iv\,\mathbb{k}$ in Fourier space. Using this we can write the redshift drift fluctuation in Fourier space as
@@
$v=\frac{1}{k^{2}}afHD_{+}\delta_{0}(\mathbb{k})\;,$ (32)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{v}{H}\right)$ $=(HafD_{+}+a^{2}Hf^{\prime}D_{+}+aHf^{2}D_{+})\frac{1}{k^{2}}\delta_{0}(\mathbb{k})$
$=\left(1+f+\frac{\mathrm{d}\ln f}{\mathrm{d}\ln a}\right)\frac{HafD_{+}}{k^{2}}\delta_{0}(\mathbb{k})\;.$ (33)
@@
$\Delta\delta z(\mathbb{k},z)=T(z)\,\delta_{0}(\mathbb{k})\frac{i}{k}\mathbb{\hat{k}}\cdot\mathbb{e}\;,$ (34)
@@
$a_{lm}^{\Delta\delta z}(z)=$ $\frac{1}{(2\pi)^{3}}\int\mathrm{d}\Omega_{\mathbb{e}}Y_{lm}^{*}(\mathbb{e})\int\mathrm{d}^{3}k\;T(z)\,\delta_{0}(\mathbb{k})\frac{i}{k}\mathbb{\hat{k}}\cdot\mathbb{e}\;\mathrm{e}^{-ir_{\mathrm{s}}\mathbb{k}\cdot\mathbb{e}}$
$=$ $-\frac{1}{(2\pi)^{3}}\int\mathrm{d}\Omega_{\mathbb{e}}Y_{lm}^{*}(\mathbb{e})\int\mathrm{d}^{3}k\;T(z)\,\delta_{0}(\mathbb{k})\frac{1}{k}\partial_{kr_{\mathrm{s}}}\mathrm{e}^{-ir_{\mathrm{s}}\mathbb{k}\cdot\mathbb{e}}\;.$ (36)
Expanding $\exp(-ir_{\mathrm{s}}\mathbb{k}\cdot\mathbb{e})$ into plane waves using the identity
$\mathrm{e}^{-ir_{\mathrm{s}}\mathbb{k}\cdot\mathbb{e}}=4\pi\sum_{l=0}^{\infty}\sum_{m=-l}^{l}(-i)^{l}j_{l}(kr)Y_{ml}(\mathbb{e})Y_{ml}^{*}(\hat{\mathbb{k}})\;,$ (37)
@@
$a_{lm}^{\Delta\delta z}(z)=-\frac{(-i)^{l}}{2(\pi)^{2}}\int\mathrm{d}^{3}k\;\frac{1}{k}T(z)\,\delta_{0}(\mathbb{k})j^{\prime}_{l}(kr_{s})Y_{ml}^{*}(\hat{\mathbb{k}})\;.$ (38)
@@
Here $P_{\delta}^{0}$ is the density fluctuation power spectrum today. In a real survey we cannot observe one precise redshift but rather consider a bin around the redshift $z\pm\mathrm{d}z$. We can account for this by adding a window function [31]
@@
[…]
As mentioned earlier, the angular power spectrum of redshift drift fluctuations has been presented in [28] in a different manner than the current presentation. The authors of [28] derived the power spectrum in a fully gauge-independent manner and also included metric fluctuations, which we neglected here. According to the results we will show in Sect. 4.1 and also the conclusions in [28], neglecting the metric perturbations is very well justified and should not influence the result significantly. Nevertheless, our results disagree with those of [28], especially for the low redshifts important in this paper. We will see that our result agrees very well with our simulation results and also with the simulation results by [13] and are therefore inclined to trust it. A full discussion of the differences between our results and those of [28] can be found in Appx. A.

### v2

@@ Abstract
The cosmological redshift drift promises to be the first observable directly measuring the evolution of the cosmic expansion rate and should be detectable with upcoming surveys by the Square Kilometre Array and the Extremely Large Telescope. To prepare for these upcoming measurements we study the redshift drift in detail using the relativistic N-body code gevolution, focusing on inhomogeneity-induced fluctuations. Using a ray-tracer, we calculate the redshift drift directly from the light cone at two different time steps. To investigate observer-dependent biases we consider 10 different observers. We find that inhomogeneity-induced fluctuations in the redshift drift can in extreme cases be of the same order as the cosmic signal for $z\lesssim 0.15$. By comparing our results to first-order perturbation theory, we find that the extreme outliers are due to peculiar motion in over-densities and can be described by first-order perturbation theory to percent precision. We calculate angular power spectra that fit very well with our predictions based on perturbation theory at linear scales and show a surprisingly large non-linear signal. This shows that redshift drift not only has the power to measure the background expansion, but could also deliver information about the velocity and acceleration fields in clusters.
@@ 1 Introduction
One of the primary tasks in cosmology is to measure the cosmic expansion rate. Such measurements constrain the spacetime geometry, the energy content of our universe, and the theory of gravity describing its evolution. Measurements of the expansion rate through supernovae of type Ia gave us the first hint that our universe is expanding in an accelerated manner, thereby prompting us to include dark energy in our model of the Universe [Riess1998; Perlmutter1999]. However, so far, most of our measurements of the expansion rate have been indirect. Measurements of supernovae for example constrain the distance-redshift relation and therewith an integrated measure of the expansion rate. A direct measurement of the cosmic expansion rate and its evolution would be a huge step forward in constraining and understanding our universe’s evolution and the phenomenology of dark energy.
Such a direct measurement of the cosmic expansion and its evolution is for the first time becoming feasible with the current generation of telescopes. Telescopes such as the Square Kilometre Array (SKA) and the Extremely Large Telescope (ELT) are predicted to be able to measure the real-time variation of the cosmological redshift [Liske2008; Kloeckner2015; Cooke2020; Rocha2023; Trost2025], the so-called redshift drift [Sandage1962; McVittie1962]. The redshift drift is the variation of the redshift, $z$, of a given source observed at two different times $\delta t_{\mathrm{o}}$ apart, i.e.
@@
The real universe is, however, only homogeneous and isotropic, i.e. described by an FLRW metric, if viewed on the largest scales. On smaller scales, the Universe is full of structures. These structures affect any cosmological observables, including the redshift drift. Indeed, the effects of structures have been identified as one of the main obstacles for obtaining unbiased redshift drift measurements [Liske2008; Cooke2020]. It was estimated in [Liske2008; Cooke2020], that the noise will be subdominant to the signal expected for observations with the ELT unless based on observations of a small number of objects with large peculiar velocity and acceleration. Furthermore, observations with the SKA will average over billions of galaxies [Kloeckner2015], meaning that peculiar motion should only introduce random noise and not bias the result. It remains important to precisely quantify the size of the noise in a realistic simulation. This is both important in order to confirm that the measurements will indeed be unbiased and do not suffer from observer-dependent effects. Furthermore, in addition to being a source of noise in the background signal, the peculiar motion of objects might provide a useful signal in itself. For example [Amendola2008] suggested that the peculiar acceleration of objects in nearby clusters could be used to map out their gravitational potential. It is therefore interesting to study whether such signals can be extracted from redshift drift measurements. The time variation of the redshift, in particular due to peculiar motion, could also be important for future gravitational wave detectors such as LISA or DECIGO [Bonvin2017; Seto2001].
[…]
@@
The paper is structured as follows. We start in Sect. 2 by introducing our simulation setup and explaining how we calculate the redshift drift. Next, we derive the redshift drift and the redshift drift fluctuation power spectrum in a perturbed FLRW universe in Sect. 3. In Sect. 4 we present the results of our simulation and compare them with the results from Sect. 3 and other results from the literature. In Sect. 5 we discuss how our results relate to current and future observational efforts, before concluding in Sect. 6.
@@ 2 Simulation Setup
The results presented in this paper are produced with the relativistic N-body code gevolution [^3] https://github.com/gevolution-code[Adamek2016_Nat]. gevolution employs a particle-mesh approach and implements the perturbed FLRW metric in the Poisson gauge
@@
where $\tau$ is the conformal time, $a$ the background scale factor and $x^{i}$ are co-moving coordinates. Here, $\Psi$, $\Phi$ are the two scalar perturbations, known as the Bardeen potentials, and $B_{i}$, $h_{ij}$ are the vector and tensor perturbations. The metric components are stored on a regular Cartesian grid and evolved according to the field equations of general relativity expanded in a weak field expansion. For the exact details of the expansion and an overview of the terms that are kept, the reader is referred to [Adamek2016]. The positions and velocities of the particles are updated by solving the geodesic equation.
In addition to the metric on the Cartesian grid and full particle positions and velocities, gevolution can output particle positions and metric fluctuations on background light cones for an observer at a given position and observation time. In this case, the metric is stored on a HEALPix [HEALPix] grid.
One can post-process the output with a raytracer to compute the redshift and position of the individual particles on the sky. We use the raytracer first introduced in [Adamek2019]. The raytracer uses an iterative procedure, shooting light rays from the observer positions towards the source, solving the geodesic equation along the ray using the perturbed metric, though only retaining the scalar perturbations. If the ray misses the source, the initial shooting angle is adjusted for the next pass until the correct geodesic is calculated.
@@
With this setup, we have chosen observers co-moving with the expansion of the universe. In a realistic scenario, the observer will be moving with respect to the background. We neglect such motion here as we are interested in the effects of the large scale structure of the universe. The contribution to the redshift drift signal from our local motion has been studied earlier [Bolejko2019; Inoue2020] and we expect its contributions to the observed redshift drift signal can be removed.
@@
[…]
@@
3 Analytical Predictions
The redshift drift in a perturbed FLRW universe was first discussed in [Linder2010; Kim2015; Marcori2018] and later derived in [Bessa2023] in a gauge-independent manner. In this paper, we will stick closest to the result by [Marcori2018], but extend it by a term that appears when observing on the light cone. We repeat the main steps of the calculation to make this paper self-contained and introduce our notation. We will only consider the dominant corrections to the redshift and redshift drift which come from the peculiar motion and not the metric fluctuations. In the derivations below we therefore assume that the wave vector is that of an FLRW spacetime and that the only redshift corrections from inhomogeneities are those from the peculiar motion of the source and observer which will be taken into account at the linear level. We thus assume that the light paths including the null tangent vector are in accordance with an FLRW spacetime with line-element
@@ 3.1 Redshift Drift
Within the geometric optics approximation the wave vector has to be $k^{\mu}=\omega\;(1,\hat{k}^{i})$ [Gravitation], where $\omega$ is the angular frequency measured by a local Lorentz observer and $\hat{k}$ is the unit direction vector pointing in the direction of propagation of the wave. Considering an observer with 4-velocity $u^{\mu}=(1,u^{i})$ one finds
@@
Since $\mathbb{n}$ is constant along the geodesic, we are free to choose it at any point along the light ray. Since we further replaced $\mathbb{n}=-\mathbb{e}$ we are now free to choose $\mathbb{e}$ at any point. We will always choose $\mathbb{e}_{\mathrm{o}}$ later, but simply keep writing $\mathbb{e}$ for ease of notation. The result (19) now agrees with the result in [Marcori2018], but is not yet the one we observe on the light cone. The second effect we have to account for comes about from the following: So far all quantities are evaluated at the background time $t$, related to hypersurfaces of constant background redshift $\bar{z}$. Observationally we do not have access to $\bar{z}$, only to the observed redshift $z=\bar{z}+\Delta z$. For any function we can relate
@@
[Bonvin2011; Bessa2023]. Within perturbation theory, such corrections need only be applied to background quantities, i.e. in our case only $H_{\mathrm{s}}$. We have
@@
which will come in handy when deriving the angular power spectrum in the next subsection. Eq. (23) agrees with the expression obtained in [Bessa2023], except for the observer-dependent term that was neglected there, and minus all terms arising from metric fluctuations, which we neglect here.
@@ 3.2 Angular Power Spectrum
We note that the power spectrum of redshift drift fluctuations has already been calculated to first order in perturbation theory in [Bessa2023] in a fully gauge-independent manner, accounting not only for the velocity perturbations but also for metric perturbations. Since we choose a different convention for $\Delta\delta z$ ([Bessa2023] neglect the $H_{0}$ term, which we retain), we present our own derivation here.
@@
$a_{lm}^{\Delta\delta z}(z)=\frac{1}{(2\pi)^{3}}\int\mathrm{d}\Omega_{\mathbb{e}}Y_{lm}^{*}(\mathbb{e})\int\mathrm{d}^{3}k\;\Delta\delta z(\mathbb{k},z)\,\mathrm{e}^{-ir\mathbb{k}\cdot\mathbb{e}}\;,$ (29)
where $r$ is the co-moving distance to our source. Introducing the velocity potential $v$ such that $\mathbb{v}=-\mathbb{\nabla}v$, we simply have $\mathbb{v}=iv\,\mathbb{k}$ in Fourier space. Using this we can write the redshift drift fluctuation in Fourier space as
@@
$v=-\frac{1}{k^{2}}afHD_{+}\delta_{0}(\mathbb{k})\;,$ (32)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{v}{H}\right)$ $=-(HafD_{+}+a^{2}Hf^{\prime}D_{+}+aHf^{2}D_{+})\frac{1}{k^{2}}\delta_{0}(\mathbb{k})$
$=-\left(1+f+\frac{\mathrm{d}\ln f}{\mathrm{d}\ln a}\right)\frac{HafD_{+}}{k^{2}}\delta_{0}(\mathbb{k})\;.$ (33)
@@
$\Delta\delta z(\mathbb{k},z)=-T(z)\,\delta_{0}(\mathbb{k})\frac{i}{k}\mathbb{\hat{k}}\cdot\mathbb{e}\;,$ (34)
@@
$a_{lm}^{\Delta\delta z}(z)=$ $-\frac{1}{(2\pi)^{3}}\int\mathrm{d}\Omega_{\mathbb{e}}Y_{lm}^{*}(\mathbb{e})\int\mathrm{d}^{3}k\;T(z)\,\delta_{0}(\mathbb{k})\frac{i}{k}\mathbb{\hat{k}}\cdot\mathbb{e}\;\mathrm{e}^{-ir\mathbb{k}\cdot\mathbb{e}}$
$=$ $\frac{1}{(2\pi)^{3}}\int\mathrm{d}\Omega_{\mathbb{e}}Y_{lm}^{*}(\mathbb{e})\int\mathrm{d}^{3}k\;T(z)\,\delta_{0}(\mathbb{k})\frac{1}{k}\partial_{kr}\mathrm{e}^{-ir\mathbb{k}\cdot\mathbb{e}}\;.$ (36)
Expanding $\exp(-ir\mathbb{k}\cdot\mathbb{e})$ into plane waves using the identity
[…]
@@
[…]
@@
Here $P_{\delta}^{0}$ is the density fluctuation power spectrum today. In a real survey we cannot observe one precise redshift but rather consider a bin around the redshift $z\pm\mathrm{d}z$. We can account for this by adding a window function [CLASSgal]
@@
To evaluate the expression (41), we use CLASS to calculate the linear matter power spectrum today $P_{\delta}^{0}(k)$ with the same cosmological parameters listed in Sect. 2 and use FFTlog-and-beyond [^6] https://github.com/xfangcosmo/FFTLog-and-beyond[Fang_2020_FFTlog] a library specifically designed to handle integrals over spherical Bessel functions and their derivatives. We use FFTlog-and-beyond to implement the first integral over the redshift $z$, transforming it to an integral over the co-moving distance $r$ to do so, and then use the composite Simpson’s rule to evaluate the remaining integral over $k$. Our implementation can be found here .
As mentioned earlier, the angular power spectrum of redshift drift fluctuations has been presented in [Bessa2023] in a different manner than the current presentation. The authors of [Bessa2023] derived the power spectrum in a fully gauge-independent manner and also included metric fluctuations, which we neglected here. According to the results we will show in Sect. 4.1 and also the conclusions in [Bessa2023], neglecting the metric perturbations is very well justified and should not influence the result significantly. Nevertheless, our results disagree with those of [Bessa2023], especially for the low redshifts important in this paper. We will see that our result agrees very well with our simulation results and also with the simulation results by [Koksbang2024] and are therefore inclined to trust it. A full discussion of the differences between our results and those of [Bessa2023] can be found in Appx. A.